# Co-STAR: Collaborative Curriculum Self-Training with Adaptive Regularization for Source-Free Video Domain Adaptation

### v1

@@ 4 Experiments
Lower bound (ResNet50) 26.5 40.0 33.5 35.1 54.1 63.6 22.8 26.3 46.2 15.4 16.8 20.7 33.4
Lower bound (ViT-B/14) 40.9 48.8 46.0 45.9 59.2 74.9 33.3 43.5 54.5 45.8 36.3 44.7 47.8
@@
Upper bound (ResNet50) 25.8 66.3 57.3 25.8 66.3 88.4 25.8 57.3 88.4 66.3 57.3 88.4 59.4
Upper bound (ViT-B/14) 58.3 82.9 70.3 58.3 82.9 97.5 58.3 70.3 97.5 82.9 70.3 97.5 77.3
Table 1: Performance on Daily-DA dataset – Symbols ‘H’, ‘K’, ‘A’, and ‘M’ represent datasets HMDB51, Kinetics, ActivityNet, and MIT/Mini-Kinetics, respectively. LB is Lower Bound and UB is Upper bound. Best results are in Bold, and the second best is underlined. For consistency, Co-STAR results with the ResNet50 backbone are compared against other benchmarks, while Co-STAR with the ViT-B/14 backbone is presented separately.
@@
Baselines. In our experiments, we compare our approach with various image-based SFUDA methods, including SFDA [21], SHOT [26], SHOT++ [28], MA [24], BAIT [52], and CPGA [35], which we adapted for evaluation on video data, as well as video-specific SFUVDA methods such as ATCoN [50], EXTERN [51], STHC [23], and DALL-V [53].
@@
For network architecture, we utilize CLIP’s visual encoders (either ResNet-50 [15] or ViT-B/14 [8]) for frame-level feature extraction, followed by a randomly initialized TRN [56] for both teacher and student models. For training the student model, we use the AdamW optimizer [29] with a learning rate of 0.001 and a weight decay of 0.2. The teacher model is updated using an EMA with a decay rate of $\delta=0.999$ (Eq. 5). The temperature parameter (Eq. 3) is set empirically at $\tau=2.0$, across all datasets. Similar to DALL-V [53], we train our model for 30 epochs in all experiments.
To generate zero-shot predictions, we leverage CLIP (ViT-B/14) [36]. Action class names are embedded using text prompts (e.g., Look, the human is {class}), and similarities between video and text embeddings are computed. These similarity scores are averaged across multiple text prompts to obtain class distribution probabilities. To reduce computational costs, these probabilities are precomputed before training. Additional implementation details are provided in the Appendix.
@@
For reference, we also report both lower and upper bounds for our Co-STAR framework using these backbones. The lower bound reflects the performance of our source-trained model on target data without adaptation, while the upper bound represents the ideal performance achievable through fully supervised training on target data. We also include the lower and upper bound accuracies (shown in the tables as LB and UB) of DALL-V [53] to provide additional context.
@@
On the Sports-DA benchmark (Table 2), Co-STAR with the ResNet50 backbone again sets a new state-of-the-art with an average accuracy of $85.6\%$, surpassing the previous best method, EXTERN, by $2.4\%$. Using the ViT-B/14 backbone further improves performance significantly, achieving average accuracy $92.4\%$.
For UCF-HMDBfull (Table 3), our method with the ResNet50 backbone performs competitively, reaching an average accuracy of 90.0%, comparable to leading methods such as STHC ($91.5\%$) and DALL-V ($91.0\%$). The ViT-B/14 variant further elevates Co-STAR’s performance, achieving an average of $92.6\%$ and establishing a new state-of-the-art on this benchmark.
@@
Lower bound (ResNet50) 83.3 74.3 83.0 70.4 46.7 44.3 67.0
Lower bound (ViT-B/14) 97.7 88.9 99.0 89.0 86.4 87.5 91.4
@@
Upper bound (ResNet50) 92.0 91.5 92.0 88.0 88.0 91.5 90.5
Upper bound (ViT-B/14) 99.4 97.3 99.4 97.4 97.4 97.3 97.9
Table 2: Performance on Sports-DA– Symbols ‘K’, ‘U’, and ‘S’ represent datasets Kinetics, UCF101, and Sports-DA, respectively. LB is Lower Bound and UB is Upper bound. Best results are in Bold, and the second best is underlined. For consistency, Co-STAR results with the ResNet50 backbone are compared against other benchmarks, while Co-STAR with the ViT-B/14 backbone is presented separately.
@@
Lower bound (ResNet50) 79.0 72.8 75.9
Lower bound (ViT-B/14) 92.1 86.8 89.4
@@
Upper bound (ResNet50) 90.5 90.5 90.5
Upper bound (ViT-B/14) 99.1 97.5 98.3
Table 3: Performance on UCF-HMDBfull - Symbols ‘H’ and ‘U’ represent datasets HMDB51 and UCF101, respectively. the datasets HMDB51, and UCF101. LB is Lower Bound and UB is Upper bound. Best results are in Bold, and the second best is underlined. For consistency, Co-STAR results with the ResNet50 backbone are compared against other benchmarks, while Co-STAR with the ViT-B/14 backbone is presented separately.
@@ 5 Ablation Study and Analysis
Impact of Curriculum Learning. Incorporating our curriculum learning strategy further enhances performance to $61.9\%$ ($+10.6\%$ over baseline), as evidenced in Table 4. This improvement validates our approach of dynamically weighting samples based on prediction reliability.
@@
Effectiveness of Pace Functions. Table 5 compares four different pace functions and a baseline (i.e. no pace function, using only reliability scores from Eq. 6) on adaptation scenarios between HMDB51 and Arid datasets in Daily-DA benchmark. The exponential pace function achieves the best performance with 48.85% average accuracy (65.0% on A→H and 32.7% on H→A), outperforming linear (46.9%), stepwise (47.4%), and sigmoid (46.6%) functions. The superior performance of exponential pacing can be attributed to two key characteristics: its gradual increase in early stages ensures the model primarily relies on the most reliable pseudo-labels when the adaptation process is most sensitive to noise, unlike linear and stepwise functions that might introduce less reliable samples too quickly. In addition, the accelerated inclusion of samples in later stages, when the model has developed more robust features, allows for more aggressive learning. This is in contrast to sigmoid’s symmetric structure, which levels off in later stages and may under-utilize the model’s improved capacity for learning from less reliable samples. Detailed formulations of all pace functions are in Appendix.
@@ 6 Limitations and Future Work
The promising results of Co-STAR, is somewhat dependent on the zero-shot capability of CLIP, which may reduce effectiveness for action classes or scenarios that are under-represented in CLIP’s training data. This reliance can limit the consistency of our collaborative pseudo-labeling strategy in such cases. Our evaluation also focuses on action recognition alone, but the method could be evaluated on other video understanding tasks with large domain gaps, such as action quality assessment [6], temporal action localization [46], and video object segmentation [12]. These tasks involve subtle cross-domain differences and could further reveal Co-STAR’s generalizability. Future work will aim to address these limitations.
@@ Appendix B Collabrative pseudo-labeling with MatchOrConf
$\footnotesize\tilde{y}=\begin{cases}\mathcal{M}(p_{\theta_{s}})&\text{if }\xi,\\ \mathcal{M}(p_{\theta_{s}})&\text{if }!\xi\text{ and }C_{s}\geq\psi_{s}\text{ and }C_{c}<\psi_{c},\\ \mathcal{M}(p_{\theta_{c}})&\text{if }!\xi\text{ and }C_{c}\geq\psi_{c}\text{ and }C_{s}<\psi_{s},\\ \mathcal{M}(p_{\theta_{s}})&\text{if }!\xi\text{ and }C_{s}\geq\psi_{s}\text{ and }C_{c}\geq\psi_{c}\text{ and }C_{s}>C_{c},\\ \mathcal{M}(p_{\theta_{c}})&\text{if }!\xi\text{ and }C_{s}\geq\psi_{s}\text{ and }C_{c}\geq\psi_{c}\text{ and }C_{c}\geq C_{s},\\ -1&\text{otherwise.}\end{cases}$ (13)
@@ Appendix C Detailed Formulations of Pace Functions
where $r_{\text{eqv}}=r(w_{\text{max}}-r)$ and the fraction of completed training scaled to $[-6,6]$ for optimal sigmoid spread.
@@
Appendix D Hyperparameters

### v2

@@ 4 Experiments
Co-STAR LB (ResNet50) 26.5 40.0 33.5 35.1 54.1 63.6 22.8 26.3 46.2 15.4 16.8 20.7 33.4
Co-STAR LB (ViT-B/14) 40.9 48.8 46.0 45.9 59.2 74.9 33.3 43.5 54.5 45.8 36.3 44.7 47.8
@@
Co-STAR UB (ResNet50) 25.8 66.3 57.3 25.8 66.3 88.4 25.8 57.3 88.4 66.3 57.3 88.4 59.4
Co-STAR UB (ViT-B/14) 58.3 82.9 70.3 58.3 82.9 97.5 58.3 70.3 97.5 82.9 70.3 97.5 77.3
Closed Gap (CG, Average %) across tasks
DALL-V [53] (ResNet50) 47.5 56.7 92.6 58.5 40.0 66.8 72.3 83.5 71.4 80.2 67.0
Co-STAR (ResNet50) 60.1 58.0 68.9 60.5 100.0 70.0 73.2 77.8 72.1 84.3 72.5
Table 1: Performance on Daily-DA dataset – Symbols ‘H’, ‘K’, ‘A’, and ‘M’ represent HMDB51, Kinetics, ARID, and MIT datasets. LB = Lower Bound, UB = Upper Bound. Best results are in Bold, and the second best is underlined. Closed Gap (CG, %) is also reported, quantifying how much of the LB–UB gap is closed (capped at 100%). Entries marked with “–” indicate cases where LB exceeds UB, making CG undefined.
@@
For network architecture, we utilize CLIP’s visual encoders (either ResNet50 [15] or ViT-B/14 [8]) for frame-level feature extraction, followed by a randomly initialized TRN [56] for both teacher and student models. For training the student model, we use the AdamW optimizer [29] with a learning rate of 0.001 and a weight decay of 0.2. The teacher model is updated using an EMA with a decay rate of $\delta=0.999$ (Eq. 5). The temperature parameter (Eq. 3) is set empirically at $\tau=2.0$, across all datasets. Similar to DALL-V [53], we train our model for 30 epochs in all experiments. Additional details, including zero-shot prediction with CLIP, are provided in the Supplementary Materials.
@@
We report both lower and upper bounds (LB, UB) for Co-STAR (and DALL-V [53] for context), where LB is the source-only performance on target data and UB is the fully supervised target performance. Alongside raw accuracies, we report the Closed Gap (CG) metric [yang2021st3d, hu2023density] for fair comparison with methods such as DALL-V that provide LB and UB. CG quantifies how much of the LB–UB gap a method closes, computed as $\text{CG}=\min\left(100,\;\frac{\text{Method}-\text{LB}}{\text{UB}-\text{LB}}\times 100\right)$. A value of 100 means the method reaches or surpasses the UB, while 0 means no improvement over LB.
@@
[…]
On the Sports-DA benchmark (Table 2), Co-STAR with the ResNet50 backbone again sets a new state-of-the-art with an average accuracy of $85.6\%$, surpassing the previous best method, EXTERN, by $2.4\%$. Using the ViT-B/14 backbone further improves performance significantly, achieving average accuracy $92.4\%$. With CG, Co-STAR obtains 79.2%, significantly higher than DALL-V’s 59.4%, highlighting significant improvement over the LB baseline.
For UCF-HMDBfull (Table 3), Co-STAR (ResNet50) performs competitively, reaching 90.0% average accuracy, comparable to leading methods such as STHC (91.5%) and DALL-V (91.0%). With CG, Co-STAR achieves 95.7%, slightly higher than DALL-V’s 95.5%. Our ViT-B/14 variant further elevates Co-STAR’s performance, achieving an average of 92.6% and establishing a new state of the art on this benchmark.
@@
Co-STAR LB (ResNet50) 83.3 74.3 83.0 70.4 46.7 44.3 67.0
Co-STAR LB (ViT-B/14) 97.7 88.9 99.0 89.0 86.4 87.5 91.4
@@
Co-STAR UB (ResNet50) 92.0 91.5 92.0 88.0 88.0 91.5 90.5
Co-STAR UB (ViT-B/14) 99.4 97.3 99.4 97.4 97.4 97.3 97.9
Closed Gap (CG, Average %) across tasks
DALL-V [53] (ResNet50) 54.0 19.8 64.4 67.6 83.5 66.9 59.4
Co-STAR (ResNet50) 90.8 36.6 100.0 80.7 91.5 75.4 79.2
Table 2: Performance on Sports-DA – Symbols ‘K’, ‘U’, and ‘S’ represent Kinetics, UCF101, and Sports-1M datasets. LB = Lower Bound, UB = Upper Bound. Best results are in Bold, and the second best is underlined. Closed Gap (CG, %) is also reported, quantifying how much of the LB–UB performance gap is closed (capped at 100%).
@@
Co-STAR LB (ResNet50) 79.0 72.8 75.9
Co-STAR LB (ViT-B/14) 92.1 86.8 89.4
@@
Co-STAR UB (ResNet50) 90.5 90.5 90.5
Co-STAR UB (ViT-B/14) 99.1 97.5 98.3
Closed Gap (CG, Average %) across tasks
DALL-V [53] (ResNet50) 100.0 91.0 95.5
Co-STAR (ResNet50) 96.5 94.9 95.7
Table 3: Performance on UCF-HMDBfull – Symbols ‘H’ and ‘U’ represent HMDB51 and UCF101, respectively. LB = Lower Bound, UB = Upper Bound. Best results are in Bold, and the second best is underlined. Closed Gap (CG, %) is also reported, quantifying how much of the LB–UB performance gap is closed (capped at 100%).
@@ 5 Ablation Study and Analysis
Impact of Curriculum Learning. Incorporating our curriculum learning strategy further enhances performance to $61.9\%$ ($10.6\%$ over baseline), as evidenced in Table 4. This improvement validates our approach of dynamically weighting samples based on prediction reliability.
@@
[…]
Effectiveness of Pace Functions. Table 5 compares four different pace functions and a baseline (i.e. no pace function, using only reliability scores from Eq. 6) on adaptation scenarios between HMDB51 and Arid datasets in Daily-DA benchmark. The exponential pace function achieves the best performance with 48.85% average accuracy (65.0% on A→H and 32.7% on H→A), outperforming linear (46.9%), stepwise (47.4%), and sigmoid (46.6%) functions. The superior performance of exponential pacing can be attributed to two key characteristics: its gradual increase in early stages ensures the model primarily relies on the most reliable pseudo-labels when the adaptation process is most sensitive to noise, unlike linear and stepwise functions that might introduce less reliable samples too quickly. In addition, the accelerated inclusion of samples in later stages, when the model has developed more robust features, allows for more aggressive learning. This is in contrast to sigmoid’s symmetric structure, which levels off in later stages and may under-utilize the model’s improved capacity for learning from less reliable samples. Detailed formulations of all pace functions are in Supplementary Materials.
@@ 6 Limitations and Future Work
The promising results of Co-STAR, are somewhat dependent on the zero-shot capability of CLIP, which may reduce effectiveness for action classes or scenarios that are under-represented in CLIP’s training data. This reliance can limit the consistency of our collaborative pseudo-labeling strategy in such cases. Our evaluation also focuses on action recognition alone, but the method could be evaluated on other video understanding tasks with large domain gaps, such as action quality assessment [6], temporal action localization [46], and video object segmentation [12]. These tasks involve subtle cross-domain differences and could further reveal Co-STAR’s generalizability. Future work will aim to address these limitations.
@@ Appendix B Collabrative pseudo-labeling with MatchOrConf
[…]
@@ Appendix C Detailed Formulations of Pace Functions
where $r_{\text{eqv}}=r(w_{\text{max}}-r)$ and the fraction of completed training scaled to [6,6][-6,6] for optimal sigmoid spread.
@@
Appendix D Impact of History Buffer Size
[…]
Appendix E Hyperparameters